### v1

# Observation of modulation-induced Feshbach resonance

Abstract
In this work, we observe a novel resonant mechanism, namely the modulation-induced Feshbach resonance. By applying a far-detuned laser to the cesium $D_{2}$ transition with intensity modulation, we periodically shake the energy levels of atomic collisional states. This periodic shaking connects the free-scattering states to shallow molecular states. At specific frequencies, we observe significant atom loss, which corresponds to the resonant coupling between these two types of states. This precisely corresponds to a form of Feshbach resonance, yet in the frequency domain rather than the magnetic-field domain. Using this method, we can directly scan the energy spectrum of molecular bound states without synthesizing any molecules. In addition to these bound states, we can also probe the molecular states embedded in the continuum, which are typically very difficult to detect by the conventional methods based on molecular synthesis. Moreover, by using a far-detuned laser instead of a magnetic field coil, it enables spatially dependent control over atomic interactions, coupling multiple levels simultaneously, and inducing new Feshbach resonances for those atoms that do not have conventional magnetic resonances. Therefore, we believe that this new resonant mechanism offers new opportunities for controlling atomic and molecular interactions in quantum simulations.
The Feshbach resonance [1] is a powerful tool in quantum many-body physics. By tuning the magnetic field, two energy levels of atomic or molecular collisional states can cross and couple with each other, leading to the divergence of scattering lengths. This resonant divergence provides a method to adjust the scattering length of ultracold atoms or molecules. Moreover, numerous new quantum phenomena have emerged through the utilization of quantum control techniques based on the Feshbach resonance [2, 3]. Following this logic, the quantum control community always demands a more versatile tool for adjusting collisional properties, as it may enable the creation of more novel quantum matters [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. The conventional Feshbach resonance [1] is based on the adjustment of the magnetic field. However, it is sometimes challenging to locally address the field or rapidly switch the field direction or magnitude. This is because the magnetic field is created by current coils, which are on a centimeter- or millimeter-scale and have a large inductance that blocks AC oscillations [12, 13, 14]. Beyond the conventional method, several tools have been developed to control the Feshbach resonance, such as the optical Feshbach resonance [22, 23, 24], electric-field-linked resonance [25, 26, 27], and so on. Each method has unique advantages in controlling atomic and molecular interactions and creating numerous new matters. Recently, a new type of Feshbach resonance induced by magnetic modulation [28, 29] has been proposed.
Following this novel concept, we first theoretically extend this idea to a more generalized scenario. Subsequently, we demonstrate this new phenomenon, the modulation-induced Feshbach resonance, by employing a far-detuned laser instead of a magnetic field coil. In previous studies [28, 29], this new resonance was based on the modulation of the scattering length. According to the Floquet theory, an AC-oscillating scattering length can enhance the DC interaction strength. By tuning the modulation frequency, it is anticipated that the effective scattering length will also diverge with respect to the frequency, which is highly similar to the conventional Feshbach resonance. A prerequisite for this modulated scattering length is the same as that of the conventional resonance: two energy levels must cross at a particular magnetic field [2, 28, 29], and any change around this specific resonant field will lead to a change in the scattering length. However, we have discovered that the pre-existence of a two-level crossing or a Feshbach resonance is not a mandatory requirement. The two energy levels of the collisional states do not need to cross with each other. All we need to do is applying a small periodic shaking to one of the energy levels, such that its energy oscillates slightly around its original value. Then, the Floquet theory can induce a new resonance between these two levels, which will directly result in the same behavior as the conventional Feshbach resonance. This provides a new way to tune the scattering length of atoms that do not have Feshbach resonances in the applicable magnetic fields.
Figure 1: (a) Experimental setup. We prepare a Bose-Einstein condensate with approximately $10^{5}$ cesium atoms in the hyperfine state of $|F=3,m_{F}=3\rangle$. The condensate is trapped by a light sheet plus a cross-dipole trap in the $x$-$y$ plane, and the magnetic field is applied along the $z$ axis. A 23-GHz-detuned light is sent along the $z$ axis with a left-hand circular polarization as depicted in panel (b). This light has an intensity modulation with a modulation frequency $\omega$, which oscillates the differential energy shift between atomic collisional states and couples the free-scattering states to the molecular states such as $4g(4)$ in panel (c).
In addition to the theoretical expansions, we conduct experimental demonstrations of this phenomenon using cesium atoms. Here, the energy level shaking is induced by the intensity modulation of a far-detuned laser, which oscillates the energy difference between atomic collisional states. Specifically, we illustrate that the free-scattering states of two atoms in the $|F=3,m_{F}=3\rangle$ state can exhibit frequency-based resonances with the molecular states $4g(4)$, $4d$ and $6s$ and a near-frequency-independent resonance with $6g(6)$. These molecular states are characterized by the quantum numbers $fl(m_{f})$ [30], where $f$ denotes the resultant of the total atomic angular momentum of the individual atoms, $l$ is the orbital angular momentum of the relative motion, and $m_{f}$ represents the projection of $f$ along the quantization axis. For states with $m_{f}=f$, we omit $(m_{f})$ for brevity. This discovery leads to two novel advantages. First, we can directly scan out most of the shallow molecular states of cesium atoms, a result that will be presented in the subsequent sections of this manuscript, including those states in the continuum. Second, in principle, we can simultaneously apply multiple frequencies to modulate the laser intensity, enabling us to bring together many energy levels (more than two) to induce this resonance. Although this might seem conceptually straightforward, it is comparable to the scenarios of two-level and three-level systems in quantum optics. A three-level system is far more complex than a two-level one, as electromagnetic-induced transparency can only occur in a three-level (or multi-level) system. Moreover, numerous quantum applications [31, 32, 33, 34, 35, 36, 37] are based on this three-level structure. Therefore, the same reasoning applies to the Feshbach resonance. A recent observation of electromagnetic-induced loss suppression [31] under the Feshbach resonance and associated atomic collisions serves as supporting evidence, suggesting the potential for tunable interactions involving more levels.
Figure 2: Typical atom loss signals in the frequency domain and dependence of the resonance position shift on the average peak intensity. In (a)-(c), we show the relative atom number $N_{m}/N_{0}$ versus the modulation frequency $\omega$ after 5 ms exposure to 0.86 W/cm2 (average peak intensity) modulation light at 18.59 G, 18.66 G and 18.70 G, respectively. We also label about the resonant peaks with corresponding molecular states $66g$ 6 and $6s$ in panels (a) and (c), while these two molecular states are strongly mixed in panel (b). The experimental atom loss signal (dark solid circle) in each panel is fitted with a Fano profile (red solid lines), and all error bars represents one standard deviation of measurements. In panel (d), we present a linear fitting example of the resonance position shift caused by the DC component of the light intensity for $4g(4)$, $4d$ molecular states and one of the $6s$-$6g(6)$ mixed molecular states at respective magnetic field of 17.27 G, 47.36 G and 18.66 G, which offers a route to compensate the resonance frequency shift.
Now we demonstrate how we extend the idea of the modulated magnetic Feshbach resonance using a two-level model. This approach retains the essential physics while neglecting microscopic details of the scattering interactions. Assume we have two states, $|\alpha\rangle$ and $|\beta\rangle$, with corresponding uncoupled energies $\hbar\omega_{\alpha}$ and $\hbar\omega_{\beta}$, where $\hbar$ is the reduced Planck constant, and define $\omega_{b}=\omega_{\alpha}-\omega_{\beta}$ to denote the energy difference. There exists a non-zero off-diagonal coupling term $\hbar\Omega/2$ between these two states due to non-zero wave function overlap. Without loss of generality, we consider that the energy of $|\beta\rangle$ is periodically shaken in the form of $\hbar\omega_{\beta}+\hbar A\cos(\omega t)$, where $A$ is the shaking amplitude and $\omega$ is the shaking frequency. This shaking term $\hbar A\cos(\omega t)$ can stem from the differential light shift between these two states caused by an intensity-modulated light. We can write the Hamiltonian $H_{0}(t)$ in matrix form as:
$H_{0}(t)=\hbar\left(\begin{array}[]{cc}\omega_{\alpha}&\Omega/2\\ \Omega/2&\omega_{\beta}+A\cos(\omega t)\end{array}\right).$ (1)
To transform this original Hamiltonian into a rotating frame, we apply a unitary transformation $U(t)$. The new Hamiltonian under transformation is $H_{1}(t)=U(t)H_{0}(t)U^{\dagger}(t)-i\hbar U(t)\partial U^{\dagger}(t)/\partial t$. When choose $U(t)$ in the diagonal form $U(t)=\textrm{Diag}\{1,e^{i\left[m\omega t+\frac{A}{\omega}\sin(\omega t)\right]}\}$, where $m$ is an integer, $H_{1}(t)$ becomes:
$H_{1}(t)=\hbar\left(\begin{array}[]{cc}\omega_{\alpha}&\frac{\Omega}{2}e^{-i\left[m\omega t+\frac{A}{\omega}\sin(\omega t)\right]}\\ \frac{\Omega}{2}e^{i\left[m\omega t+\frac{A}{\omega}\sin(\omega t)\right]}&\omega_{\alpha}-\omega_{b}-m\omega\end{array}\right).$ (2)
By using the expansion $e^{i\frac{A}{\omega}\sin(\omega t)}=\sum_{n}J_{n}(A/\omega)e^{in\omega t}$, where $J_{n}(x)$ is the $n$-th order Bessel function of the first kind, we rewrite $H_{1}$ as:
$\frac{H_{1}(t)}{\hbar}=\omega_{\alpha}+$
$\left(\begin{array}[]{cc}0&\sum_{n}\frac{\Omega}{2}J_{n}(\frac{A}{\omega})e^{-i(m+n)\omega t}\\ \sum_{n}\frac{\Omega}{2}J_{n}(\frac{A}{\omega})e^{i(m+n)\omega t}&-\omega_{b}-m\omega\end{array}\right).$ (5)
The first term $\omega_{\alpha}$ is a constant and can be ignored. The second term describes the interaction between the two states. By retaining the resonant term under the condition $n=-m$, we can simplify the interaction Hamiltonian to:
$\frac{H_{1}(t)}{\hbar}-\omega_{\alpha}=\left(\begin{array}[]{cc}0&\frac{\Omega}{2}(-1)^{m}J_{m}(\frac{A}{\omega})\\ \frac{\Omega}{2}(-1)^{m}J_{m}(\frac{A}{\omega})&-\omega_{b}-m\omega\end{array}\right).$ (6)
Therefore, when $m\omega$ approaches $-\omega_{b}=\omega_{\beta}-\omega_{\alpha}$, the system becomes a resonant-coupled two-level system. Detailed derivations of its effects on scattering properties and scattering lengths are provided in the Supplementary Materials [38]. Here, we present the conclusion that due to the resonant coupling from $|\alpha\rangle$ to $|\beta\rangle$, the effective scattering length $a_{s}$ of the incoming state $|\alpha\rangle$ has the form:
$a_{s}=a_{\textrm{BK}}\left(1-\frac{\Delta_{m}}{-m\omega-\omega_{0}}\right).$ (7)
Figure 3: Multiple resonances of the $4g(4)$ molecular state. (a) Sweeping the modulation frequency at 19.41 G, we observe three distinct atom loss peaks, located at 228.7 kHz, 114.3 kHz, and 75.0 kHz, corresponding to the first-, second-, and third-order resonance frequencies of the $4g(4)$ state, respectively. (b) Magnetic field scan under a fixed modulation frequency of 150 kHz, reveals an optically shifted Feshbach resonance at 19.70 G, along with the first- and second-order modulation resonances located at 19.51 G, 19.32 G (left side), 19.90 G and 20.08 G (right side), respectively. The average peak intensity of used modulation light in both two panels is 0.87 W/cm2.
Here $a_{\textrm{BK}}$ is the background scattering length, $\Delta_{m}$ is the width of the resonance and $\omega_{0}$ is close to $\omega_{b}$. It exactly corresponds to the form of Feshbach resonance, but with the magnetic field replaced by an integer multiple of the modulation frequency $\omega$, which we use the term, modulation-induced Feshbach resonance, to describe it. Comparing to the previous study by modulating the scattering length, here the two energy levels $|\alpha\rangle$ and $|\beta\rangle$ do not need any crossing or preexisted Feshbach resonance, as it only need tiny level shaking.
Figure 4: The energy spectrum of cesium molecular states. We plot the modulation frequencies $\omega$ (or energies) versus the magnetic fields $B$. The vertical axis is plotted upside down in order to provide the same illustration of the bound state plot. Here, we use filled symbols to plot data corresponding to molecular bound states whose energy is lower than 0, and we use unfilled symbols to plot data corresponding to molecular states in the continuum whose energy is higher than 0. The black solid line along $\omega=0$ corresponds to the free-scattering states. The other solid lines are the theoretical models calculated for molecular states $6g(6)$, $6s$, $4g(4)$ and $4d$. The $4g(4)$ and $4d$ data are obtained at a average peak intensity 0.86 W/cm2 with corresponding theoretical lines shifted from their original Feshbach resonance points to the optically shifted ones measured in experiments, while the $6s$ and $66g$ 6 data are obtained by varying the average peak intensity and compensating the DC component of the light shift by linear fitting.
We now turn to the experimental implementation. We prepare a Bose-Einstein condensate consisting of approximately $10^{5}$ cesium atoms, with all atoms initialized in the $6S$ hyperfine ground state $|F=3,m_{F}=3\rangle$ [38]. A magnetic field $B$ is applied along the $z$-axis (as depicted in Fig. 1a) and serves as the quantization axis. Subsequently, an intensity-modulated light with a left-hand circularly polarization and a 23 GHz red-detuning relative to the cesium $D_{2}$ transition of $F=3\rightarrow F^{\prime}=4$ (Fig. 1b), is directed along the $z$-axis towards the atoms, inducing the light shift on the free-scattering states and molecular states [38]. The average peak intensity is varied from 0.2 W/cm2 to 1.7 W/cm2 in our experiments, while the modulation depth is fixed to be $86\%$. Then we vary the modulation frequencies $\omega$ with the aim of coupling the molecular states (Fig. 1c). After irradiating the atoms with the light for 1-20 ms, we switch off the light and measure the remaining atom number [15].
In Fig. 2a-c, we present data obtained at different magnetic fields (18.59, 18.66, and 18.70 G). Here the modulation-induced resonances are exhibited by the significant atom loss signals of the relative atom number $N_{m}/N_{0}$, where $N_{m}$ is the remaining atom number after modulation, and $N_{0}$ is the atom number far away from resonance positions. These loss signals correspond to the molecular state $6s$ and $6g(6)$. Especially, the energy level of $6g(6)$ is nearly parallel to the incoming scattering state (both atoms in $|F=3,m_{F}=3\rangle$) as a function of the magnetic field, and this state does not have a direct energy crossing with the scattering state at low magnetic field regime [39]. These experimental results confirm our hypothesis that modulation can induce a resonance between two states without the pre-existence of a conventional Feshbach resonance. Furthermore, we also observe other atom loss features correspond to the molecular states $4g(4)$ and $4d$, and their resonant frequencies exhibit obvious dependence on the magnetic field like the molecular state $6s$. Notably, the measured atom loss peaks are all red-shifted due to the unequal DC component of the light shift between the free-scattering states and the molecular states. This shift can be determined and compensated by way of linearly fitting between the atom loss peak position and the average peak intensity, as shown in Fig. 2d. The atom-loss characteristics as a function of frequency precisely mirror the form of the conventional Feshbach resonance, where atom-loss features are measured as a function of the magnetic field. This type of resonance in the frequency domain is what we term the modulation-induced Feshbach resonance.
[…]
In Fig. 4, we plot the measured resonant frequencies as a function of the magnetic field. Simultaneously, we plot the theoretical values [30, 40, 39] of these molecular states for comparison (solid lines). The vertical axis of the modulation frequency $\omega$ is inverted for the bound states whose energies are below those of the free-scattering states. In addition to the bound states, we also identify the molecular states embedded in the continuum. The unfilled symbols in the plot represent our measurement results for these states in the continuum. Our experimental data (mainly shifted by the DC component of the light intensity) reveal the $6g(6)$, $6s$, $4d$, and $4g(4)$ molecular states. By comparing experimental data with theoretical calculations, we observe excellent agreement between them, demonstrating the effectiveness of our method for probing these molecular states.
Beyond the overall view of the energy spectrum, there are also intriguing detailed features observable in this plot. In Fig. 4c, we zoom in on the crossing region of the $6g(6)$ and $6s$ states, where the states are obtained by eliminating the light shift. This region exhibits the characteristic features of a Landau-Zener crossing, where two energy levels intersect and a coupling term between them creates an energy gap. By fitting the data to the Landau-Zener model where the energies are decided by $E_{\pm}=\left[{E_{i}+E_{j}\pm\sqrt{(E_{i}-E_{j})^{2}+V^{2}_{ij}}}\right]/{2}$, we determine that the Rabi frequency $V_{ij}/h$ of this coupling between these two states is 25 kHz.
In conclusion, we have experimentally observed a novel resonant mechanism, the modulation-induced Feshbach resonance, through the utilization of an intensity-modulated light. This new resonance shares a similar form with the conventional Feshbach resonance, yet it occurs in the frequency domain rather than relying on a tunable magnetic field. It enables us to directly map out the energy spectrum of shallow molecular states, including those embedded in the continuum. Given the highly versatile tunability of lasers, in principle, we can bring together more than two energy levels to induce the resonance, which we anticipate will surpass the capabilities of the conventional Feshbach resonance. We look forward to the further development of future experimental investigations and associated applications based on a three-level resonance.
This work is supported by the National Natural Science Foundation of China and the National Key Research and Development Program of China.